OpenStudy (anonymous):

How long does is take a 25% efficient 3400W crane to lift a 720kg block 44m?

OpenStudy (anonymous):

Work done W = mgh = 720*10*44 = 316800 Time t = W/P = 316800/(3400*25%) ~ 373 s :D

OpenStudy (anonymous):

thanks looks so simple know that i see it laid out there. i was so close on my own i was over thinking the problem. thanks alot

OpenStudy (anonymous):

wish i could help you with your problem

OpenStudy (anonymous):

You're welcome :D